# publications.bib

### v3

@@ Abstract.
We generalize the torsion vanishing results of [CS1, CS2, Ko, San]. Our results apply to the cohomology of general Shimura varieties $(\mathbf{G},X)$ of PEL type $A$ or $C$, localized at a suitable maximal ideal $\mathfrak{m}$ in the spherical Hecke algebra at primes $p$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a group for which we know the Fargues-Scholze local Langlands correspondence is the semi-simplification of a suitably nice local Langlands correspondence, as shown in [FS, Ham1, HKW, BMNH]. This is accomplished by combining Koshikawa’s technique [Ko], the theory of geometric Eisenstein series over the Fargues-Fontaine curve [Ham2], the work of Santos [San] describing the structure of the fibers of the minimally and toroidally compactified Hodge-Tate period morphism for general PEL type Shimura varieties of type $A$ or $C$, and ideas developed by Zhang [Zha] on comparing Hecke correspondences on the moduli stack of $G$-bundles with the cohomology of Shimura varieties. In the process, we also establish a description of the generic part of the cohomology that bears resemblance to the work of Xiao-Zhu [XZ]. Moreover, we also construct a filtration on the compactly supported cohomology that differs from Mantovan’s filtration in the case that the Shimura variety is non-compact, allowing us to circumvent some of the circumlocutions taken in [CS2, Ko]. Our method showcases a very general strategy for proving such torsion vanishing results, and should bear even more fruit once the inputs are generalized. Motivated by this, we formulate an even more general torsion vanishing conjecture (Conjecture 6.6).
@@ 1. Introduction
Let $\mathbf{G}$ be a connected reductive group over $\mathbb{Q}$ admitting a Shimura datum $(\mathbf{G},X)$, and let $\mathbb{A}$ (resp. $\mathbb{A}_{f}$) denote the adeles (resp. finite adeles) of $\mathbb{Q}$. Fix a prime number $p>0$ and let $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ be the base-change to $\mathbb{Q}_{p}$. We will assume that $G$ is unramified so that there exists a hyperspecial subgroup $K^{\mathrm{hs}}_{p}\subset G(\mathbb{Q}_{p})$ and a Borel $B$ surjecting onto a maximal torus $T$ which we now fix. We consider the open compact subgroup $K:=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, where $K^{p}\subset\mathbf{G}(\mathbb{A}_{f}^{p})$ denotes a sufficiently small level away from $p$. Let $\mathrm{Sh}(\mathbf{G},X)_{K}$ denote the corresponding Shimura variety defined over the reflex field $E$. Given a prime $p\neq\ell$, we will be interested in understanding the $\ell$-torsion cohomology groups
@@
$H_{K_{p}^{\mathrm{hs}}}:=\overline{\mathbb{F}}_{\ell}[K_{p}^{\mathrm{hs}}\backslash G(\mathbb{Q}_{p})/K_{p}^{\mathrm{hs}}]$
@@
specified by a semisimple element $\phi_{\mathfrak{m}}(\mathrm{Frob}_{\mathbb{Q}_{p}})$. In particular, if $T$ denotes the maximal torus of $G$ then $\phi_{\mathfrak{m}}$ is induced from a parameter $\phi_{\mathfrak{m}}^{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\subset\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ factoring through the $L$-group of the maximal torus. Now, recall that the irreducible representations of $\phantom{}{}^{L}T$ correspond to the $\Gamma$-orbits $\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ of geometric dominant cocharacters of $G$. We have the following definition.
@@
[Ham2, Definition 1.4] Given a toral $L$-parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$, we say that $\phi_{T}$ is generic if, for all $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ corresponding to a $\Gamma$-orbit of coroots, we have that the complex $R\Gamma(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial. Similarly, we say that $\mathfrak{m}$ is generic if $\phi_{\mathfrak{m}}^{T}$ is a generic toral parameter.
If $G=\mathrm{GL}_{n}$ then this coincides with the notion of decomposed generic considered in [CS1, Definition I.9]. We set $d=\mathrm{dim}(\mathrm{Sh}(\mathbf{G},X)_{K})$. Motivated by [CS1, Theorem 1.1] and [CS2, Theorem 1.1], we make the following conjecture.
@@
[…]
@@
[CS1, CS2, Ko, San] Assume that $(\mathbf{G},X)$ is a PEL type Shimura datum of type $A$. If the prime $p$ splits completely in $F$ then Conjecture 1.2 is true.
@@
Koshikawa proved this under the assumption that $B=F$ and $V=F^{2n}$, and the global unitary group $\mathbf{G}$ is quasi-split, as well as in the case when $p$ is split in $F$ and the Shimura variety is compact. These additional assumptions were removed in the PhD thesis of Santos [San].
@@
Caraiani-Scholze actually proved a slightly different result. More precisely, let $S$ be a set of finite places not containing $p$ such that $\mathbf{G}$ is unramified and $K^{p}$ is hyperspecial away from $S$. Consider a maximal ideal $\mathfrak{m}\subset\mathbb{T}^{S}$ in the spherical Hecke algebra such that $\mathfrak{m}$ is generic at $p$. Caraiani-Scholze show that the localization at $\mathfrak{m}^{p}\subset\mathbb{T}^{S\cup\{p\}}$ is concentrated in the relevant degrees.
@@
In the case of Harris-Taylor Shimura varieties, there is also work of Boyer [Boy], which describes the localization at non-generic maximal ideals.
@@
We believe that Conjecture 1.2 is true under the weaker hypothesis that $H^{2}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial for all $\Gamma$-orbits of coroots $\alpha$, as is shown in [CS2, San, Ko] in their particular case. However, the theory of geometric Eisenstein series which we will invoke in this paper becomes more complicated in this case (See the discussion around [Ham2, Conjecture 1.29]), and so a proof of this Theorem using our methods would require more deeply understanding geometric Eisenstein series when this assumption is dropped (cf. Remark 6.8).
@@
Caraiani-Scholze [CS1, CS2] proved their results under some small restrictions, which Koshikawa [Ko] was able to remove by using compatibility of the Fargues-Scholze local Langlands correspondence with the semi-simplification of the Harris-Taylor correspondence for $\mathrm{GL}_{n}$. In the process, Koshikawa exhibited a much more flexible method for proving Theorem 1.3. The goal of the current paper is to expand the scope of Koshikawa’s technique, motivated by work of the first author on geometric Eisenstein series in the Fargues-Fontaine setting [Ham2]. We then carry the strategy out in some particular cases using work on local-global compatibility of the Fargues-Scholze local Langlands correspondence beyond the case of $\mathrm{GL}_{n}$.
@@
One of the basic ingredients is the perspective on Mantovan’s product formula provided by the Hodge-Tate period morphism. To explain this, we let $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ denote the minuscule geometric dominant cocharacter of $G$ determined by the Hodge cocharacter of $X$ and an isomorphism $j:\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$ which we fix from now on. We consider the Kottwitz set $B(G)$ and with it the subset $B(G,\mu)\subset B(G)$ of $\mu$-admissible elements. Let $\mathfrak{p}|p$ be the prime dividing $p$ in the reflex field $E$, induced by the embedding ¯¯$\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ given by the isomorphism $j$. We let $E_{\mathfrak{p}}$ be the completion at $\mathfrak{p}$, $C:=\hat{\overline{E}}_{\mathfrak{p}}$ be the completion of the algebraic closure, and $\breve{E}_{\mathfrak{p}}$ be the compositum of $E_{\mathfrak{p}}$ with the completion of the maximal unramified extension of $\mathbb{Q}_{p}$. We recall that, attached to each element $b\in B(G,\mu)$, there exists a diamond
@@
of meromorphy $\mu$ between the $G$-bundle $\mathcal{E}_{b}$ corresponding to $b$ and the trivial $G$-bundle. This space has an action by $G(\mathbb{Q}_{p})=\mathrm{Aut}(\mathcal{E}_{0})$ and $J_{b}(\mathbb{Q}_{p})\subset\mathrm{Aut}(\mathcal{E}_{b})$, where $J_{b}$ is the $\sigma$-centralizer of $b$. This allows us to consider the quotients
@@
at hyperspecial level with torsion coefficients. This has an action of $W_{E_{\mathfrak{p}}}\times J_{b}(\mathbb{Q}_{p})\times H_{K_{p}^{\mathrm{hs}}}$. Now, the Mantovan product formula tells us that if we look at $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})$ then this should always admit a filtration in the derived category whose graded pieces are
$(R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))[2d_{b}]\otimes_{\mathcal{H}(J_{b})}R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$
@@
    $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf on $\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}}$ with trivial Weil group action and $J_{b}(\mathbb{Q}_{p})$ action as defined in [Ko, Lemma 7.4].
Such a filtration should always exist, but is not currently proven in general. In the case that the Shimura datum $(\mathbf{G},X)$ is PEL of type $A$ or $C$, a modern proof of this result can be found in [Ko, Theorem 7.1].
@@
We first discuss point (1). One of the key observations underlying Koshikawa’s method was that the cohomology of the space $\mathrm{Sht}(G,b,\mu)_{\infty}$ computes the action of a Hecke operator $T_{\mu}$ corresponding to $\mu$ on $\mathrm{Bun}_{G}$ the moduli stack of $G$-bundles of the Fargues-Fontaine curve. The Hecke operators commute with the action of the excursion algebra on $\mathrm{Bun}_{G}$, and the action of the excursion algebra on a smooth irreducible representation $\rho$, viewed as a sheaf on $\mathrm{Bun}_{G}$, determines the Fargues-Scholze parameter of $\rho$. It follows that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ as a complex of smooth $J_{b}(\mathbb{Q}_{p})$-modules will have irreducible constituents $\rho$ with Fargues-Scholze parameter $\phi_{\rho}^{\mathrm{FS}}$ equal to $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. When $\mathbf{G}_{\mathbb{Q}_{p}}=G$ is a product of $\mathrm{GL}_{n}$s as in Theorem 1.3 (by the assumption that $p$ splits in $F$), it follows from the work of Hansen-Kaletha-Weinstein [HKW, Theorem 1.0.3] that the Fargues-Scholze correspondence for $J_{b}(\mathbb{Q}_{p})$ with rational coefficients agrees with the semi-simplification of the Harris-Taylor correspondence, where we recall that $J_{b}$ is a product of inner forms of $\mathrm{GL}_{n}$s in this case. In particular, using that $\mathfrak{m}$ is generic, it follows that $\phi_{\rho}^{\mathrm{FS}}=\phi_{\mathfrak{m}}$ must lift to a $\overline{\mathbb{Z}}_{\ell}$ parameter which is also generic in the analogous sense, and the condition of generic implies that the lift cannot come from the semi-simplification of a parameter with non-trivial monodromy. Using this, one can deduce that such a $\rho$ only exists if the group $J_{b}$ is quasi-split. In this particular case ($G$ is a product of $\mathrm{GL}_{n}$s), this can only happen if $b\in B(G,\mu)$ corresponds to the ordinary element.
@@
This argument of Koshikawa was formalized and generalized further in work of the first author [Ham2]. In particular, it was noted that, for a general quasi-split $G$ and $\mathfrak{m}$ generic, the cohomology $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ will only be non-trivial if $b\in B(G,\mu)_{\mathrm{un}}:=B(G)_{\mathrm{un}}\cap B(G,\mu)$, where $B(G)_{\mathrm{un}}$ is the set of elements lying in the image of the map $B(T)\rightarrow B(G)$, assuming that the Fargues-Scholze local Langlands correspondence has certain expected properties (Assumption 4.4). These unramified elements will be precisely the elements for which $J_{b}$ is quasi-split. The set $B(G,\mu)_{\mathrm{un}}$ corresponds to Weyl group orbits of weights in the representation $V_{\mu}$ of $\hat{G}$ restricted to $\hat{G}^{\Gamma}$. In particular, if $G$ is split then, since $\mu$ is minuscule, $B(G,\mu)_{\mathrm{un}}$ consists of only one element, corresponding to the unique Weyl group orbit of the highest weight. This is the situation occurring in the previous paragraph. Moreover, the contribution of the cohomology of this shtuka space is easily understood, and the problem completely reduces to controlling the cohomology of $\mathrm{Ig}^{b}$ when $b\in B(G,\mu)_{\mathrm{un}}$ is the $\mu$-ordinary element. However, if $G$ is not split then the restriction of $V_{\mu}$ to $\hat{G}^{\Gamma}$ may have multiple Weyl group orbits of weights. In particular, one needs to control the cohomology groups

### v5

# Torsion Vanishing for some Shimura Varieties

@@ Abstract.
We generalize the torsion vanishing results of [7, 8, 37, 54]. Our results apply to the cohomology of general Shimura varieties $(\mathbf{G},X)$ of PEL type AC, localized at a suitable maximal ideal $\mathfrak{m}$ in the spherical Hecke algebra at primes $p$ such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a group for which we know the Fargues-Scholze local Langlands correspondence is the semi-simplification of a suitably nice local Langlands correspondence, as shown in [16, 27, 31, 47]. This is accomplished by combining Koshikawa’s technique [37], the theory of geometric Eisenstein series over the Fargues-Fontaine curve [24], the work of Santos [54] describing the structure of the fibers of the minimally and toroidally compactified Hodge-Tate period morphism for general PEL type Shimura varieties of type AC, and ideas developed by Zhang [68] on comparing Hecke correspondences on the moduli stack of $G$-bundles with the cohomology of Shimura varieties. In the process, we also establish a description of the generic part of the cohomology that bears resemblance to the work of Xiao-Zhu [67]. Moreover, we also construct a filtration on the compactly supported cohomology that differs from Mantovan’s filtration in the case that the Shimura variety is non-compact,allowing us to circumvent some of the circumlocutions taken in [8, 37]. Our method showcases a very general strategy for proving such torsion vanishing results, and should bear even more fruit once the inputs are generalized. Motivated by this, we formulate an even more general torsion vanishing conjecture (Conjecture 6.6).
@@ 1. Introduction
1.0.1. The Main Result
Let $\mathbf{G}$ be a connected reductive group over $\mathbb{Q}$ admitting a Shimura datum $(\mathbf{G},X)$ which we fix from now on. Fix a prime number $p>0$ and let $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ be the base-change to $\mathbb{Q}_{p}$. We will assume that $G$ is unramified so that there exists a hyperspecial subgroup $K^{\mathrm{hs}}_{p}\subset G(\mathbb{Q}_{p})$ and a Borel $B\subset G$ surjecting onto a maximal torus $T\subset G$ which we now fix. We consider the open compact subgroup $K:=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, where $K^{p}\subset\mathbf{G}(\mathbb{A}_{f}^{p})$ denotes a sufficiently small level away from $p$. Let $\mathrm{Sh}(\mathbf{G},X)_{K}$ denote the corresponding Shimura variety defined over the reflex field $E$. Given a prime $p\neq\ell$, we will be interested in understanding the $\ell$-torsion cohomology groups
@@ 1.0.1. The Main Result
[…]
@@
up to $\hat{G}-$conjugacy specified by a semisimple element $\phi_{\mathfrak{m}}(\mathrm{Frob}_{\mathbb{Q}_{p}})$. In particular, the $L$-parameter $\phi_{\mathfrak{m}}$ up to conjugacy is induced from a parameter $\phi_{\mathfrak{m}}^{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})\subset\phantom{}^{L}G(\overline{\mathbb{F}}_{\ell})$ factoring through the $L$-group of the maximal torus. Now, recall that the irreducible representations of $\phantom{}{}^{L}T$ correspond to the $\Gamma$-orbits $\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ of geometric cocharacters of $T$. We have the following definition.
@@
[24, Definition 1.2] Given a toral $L$-parameter $\phi_{T}:W_{\mathbb{Q}_{p}}\rightarrow\phantom{}^{L}T(\overline{\mathbb{F}}_{\ell})$, we say that $\phi_{T}$ is generic if, for all $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})/\Gamma$ corresponding to a $\Gamma$-orbit of coroots, we have that the complex $R\Gamma(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial. Similarly, we say that $\mathfrak{m}$ is generic if the $L$-parameter $\phi_{\mathfrak{m}}^{T}$ is a generic toral parameter, where we note that this only depends on $\phi_{\mathfrak{m}}$ up to $\hat{G}-$conjugacy.
If $G=\mathrm{GL}_{n}$ then this coincides with the notion of decomposed generic considered in [7, Definition I.9]. We set $d=\mathrm{dim}(\mathrm{Sh}(\mathbf{G},X)_{K})$. Motivated by [7, Theorem 1.1] and [8, Theorem 1.1], we make the following conjecture.
@@
We first recall the motivating situation of Caraiani-Scholze [7, 8]. Let $F/\mathbb{Q}$ be a CM field, and let $(B,*,V,\langle\cdot,\cdot\rangle)$ be a PEL datum with $B$ a central simple $F$-algebra and $V$ a non-zero finite type left $B$-module. Let $(\mathbf{G},X)$ denote the Shimura datum attached to it, where $\mathbf{G}$ is the connected reductive group over $\mathbb{Q}$ defined by the automorphisms of $V$ preserving the choice of $B$-linear pairing $\langle\cdot,\cdot\rangle$ up to similitude. We have the following result.
@@
[7, 8, 37, 54] Assume that $(\mathbf{G},X)$ is a PEL type Shimura datum of type $A$. If the prime $p$ splits completely in $F$ then Conjecture 1.2 is true.
@@
Koshikawa proved this under the assumption that $B=F$ and $V=F^{2n}$, and the global unitary group $\mathbf{G}$ is quasi-split, as well as in the case when $p$ is split in $F$ and the Shimura variety is compact. These additional assumptions were removed in the PhD thesis of Santos [54].
@@
Caraiani-Scholze actually proved a slightly different result. More precisely, let $S$ be a set of finite places not containing $p$ such that $\mathbf{G}$ is unramified and $K^{p}$ is hyperspecial away from $S$. Consider a maximal ideal $\mathfrak{m}\subset\mathbb{T}^{S}$ in the prime-to-$S$ Hecke algebra. This defines a global Galois representation $\rho_{\mathfrak{m}}$, and we can consider the restriction $\rho_{\mathfrak{m}}|_{\mathrm{Gal}(\bar{E}_{v}/E_{v})}$ which is unramified. If the associated toral $L$-parameter is generic as in Definition 1.1, then Caraiani-Scholze show (under some technical global assumptions) that the localization at $\mathfrak{m}^{p}\subset\mathbb{T}^{S\cup\{p\}}$ is concentrated in the same degrees as Conjecture 1.2.
@@
In the case of Harris-Taylor Shimura varieties, there is also work of Boyer [4], which describes the localization of the torsion cohomology at non-generic maximal ideals.
@@
We believe that Conjecture 1.2 is true under the weaker hypothesis that $H^{2}(W_{\mathbb{Q}_{p}},\alpha\circ\phi_{T})$ is trivial for all $\Gamma$-orbits of coroots $\alpha$, as is shown in [8, 54, 37] in their particular case. However, the theory of geometric Eisenstein series which we will invoke in this paper becomes more complicated in this case (See the discussion around [24, Conjecture 1.24]), and so a proof of this Theorem using our methods would require more deeply understanding geometric Eisenstein series when this assumption is dropped (cf. Remark 6.8). In certain cases however the theory of geometric Eisenstein series becomes unnecessary, and in this case an argument using compatibility (See Assumption 4.5) of Fargues-Scholze with some form of local Langlands is sufficient, as in the style of the arguments of [37, 54] (See Remark 4.28 for details).
@@
Caraiani-Scholze [7, 8] proved their results under some technical restrictions of a global nature, which Koshikawa [37] was able to remove by using compatibility of the Fargues-Scholze local Langlands correspondence with the semi-simplification of the Harris-Taylor correspondence for $\mathrm{GL}_{n}$. In the process, Koshikawa exhibited a much more flexible method for proving Theorem 1.3. The goal of the current paper is to expand the scope of Koshikawa’s technique, motivated by work of the first author on geometric Eisenstein series in the Fargues-Fontaine setting [24]. We then carry the strategy out in some particular cases using work on local-global compatibility of the Fargues-Scholze local Langlands correspondence beyond the case of $\mathrm{GL}_{n}$, as studied in [24, 47]. Our main result is the following.
Theorem 1.8.
(Theorem 5.2) Suppose $(\mathbf{G},X)$ is a PEL datum of type AC satisfying Assumption 1.13 such that $\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. Let $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ be a generic maximal ideal. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology of $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
Remark 1.9.
This notably allows one to relax the assumption in [7, 8, 37, 54] that the prime $p$ splits in $F$ in the setting of Theorem 1.3, answering a question of Caraiani.
Here is the table summarizing our local constraints:
$G$ Constraint on $G$ $\ell$ $p$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GSp}_{4})$ $L=\mathbb{Q}_{p}$
$L/\mathbb{Q}_{p}$ unramified $(\ell,[L:\mathbb{Q}_{p}]=1$ $p\neq 2$
$\mathrm{Res}_{L/\mathbb{Q}_{p}}(\mathrm{GU}_{2})$ $L/\mathbb{Q}_{p}$ unramified $(\ell,2[L:\mathbb{Q}_{p}])=1$
$G=\mathrm{U}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G=\mathrm{GU}_{n}(L/\mathbb{Q}_{p})$ $n$ odd $L$ unramified $\ell\neq 2$
$G(\mathrm{Sp}_{4,L})$ $L/\mathbb{Q}_{p}$ unramified, $L\neq\mathbb{Q}_{p}$ $(\ell,[L:\mathbb{Q}_{p}])=1$ $p\neq 2$
We can also easily deduce a result for some abelian type Shimura varieties, such as Hilbert modular varieties, from the above result, which recovers results similar to that of Caraiani-Tamiozzo [9, Theorem B] (See Corollary 5.7).
Corollary 1.10.
(Corollary 5.6) Suppose $(\mathbf{G},X)$ is an abelian-type Shimura datum which has an associated PEL-type datum $(\mathbf{G}_{1},X_{1})$ of type AC satisfying Assumption 1.13 and such that $\mathbf{G}^{\mathrm{der}}\simeq\mathbf{G}_{1}^{\mathrm{der}}$, $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$ is unramified and $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ is a product of groups as in Table (1) with $p$ and $\ell$ satisfying the corresponding conditions. We write $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ and $G_{1}:=\mathbf{G}_{1,\mathbb{Q}_{p}}$. We assume that $\mathfrak{m}\subset H_{K_{p}^{\mathrm{hs}}}$ is a generic maximal ideal. Then, for a level $K=K^{p}K_{p}^{\mathrm{hs}}\subset\mathbf{G}(\mathbb{A}_{f})$, the cohomology $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$ (resp. $R\Gamma_{c}(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})_{\mathfrak{m}}$) is concentrated in degrees $d\leq i\leq 2d$ (resp. $0\leq i\leq d$).
We now explain how we are able to establish these results, as well as several other interesting results on the perverse $t$-exactness of Hecke operators on $\mathrm{Bun}_{G}$ (Theorem 1.11) and the splitting of Mantovan’s filtration on the cohomology of the Shimura variety (Theorem 1.17) along the way.
1.0.2. Proof Sketch of the Main Theorem
@@
One of the basic ingredients used in verifying Conjecture 1.2 in all known instances is the perspective on Mantovan’s product formula provided by the Hodge-Tate period morphism. To explain this, we let $\mu\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ denote the minuscule geometric dominant cocharacter of $G$ determined by the inverse of the Hodge cocharacter of $X$ and an isomorphism $j:\mathbb{C}\simeq\overline{\mathbb{Q}}_{p}$ which we fix from now on. We consider the Kottwitz set $B(G)$, and with it the subset $B(G,\mu)\subset B(G)$ of $\mu$-admissible elements. Let $\mathfrak{p}|p$ be the prime dividing $p$ in the reflex field $E$ induced by the embedding $\overline{\mathbb{Q}}\rightarrow\overline{\mathbb{Q}}_{p}$ given by the isomorphism $j$. We let $E_{\mathfrak{p}}$ be the completion at $\mathfrak{p}$, $C:=\hat{\overline{E}}_{\mathfrak{p}}$ be the completion of the algebraic closure, and $\breve{E}_{\mathfrak{p}}$ be the compositum of $E_{\mathfrak{p}}$ with $\breve{\mathbb{Q}}_{p}$. We recall that, attached to each element $b\in B(G,\mu)$, there exists a diamond
@@
of meromorphy $\mu$ between the $G$-bundle $\mathcal{E}_{b}$ corresponding to $b$ and the trivial $G$-bundle. This space has an action by $\underline{G(\mathbb{Q}_{p})}=\mathrm{Aut}(\mathcal{E}_{0})$ and $\underline{J_{b}(\mathbb{Q}_{p})}\subset\mathrm{Aut}(\mathcal{E}_{b})$, where $J_{b}$ is the $\sigma$-centralizer of $b$. This allows us to consider the quotients
@@
at hyperspecial level with torsion coefficients. This has an action of $W_{E_{\mathfrak{p}}}\times J_{b}(\mathbb{Q}_{p})\times H_{K_{p}^{\mathrm{hs}}}$ [^1] It is clear that the space has an action of the inertia group $I_{E_{\mathfrak{p}}}$ of $E_{\mathfrak{p}}$ from the definition; however, the space $\mathrm{Sht}(G,b,\mu)_{\infty}\rightarrow\mathop{\rm Spd}(\breve{E}_{\mathfrak{p}})$ also has a canonical (non-effective) Frobenius descent datum giving a full action of $W_{E_{\mathfrak{p}}}$ on the cohomology (See [16, Section IX.3]). Now, the Mantovan product formula tells us that if we look at $R\Gamma(\mathrm{Sh}(\mathbf{G},X)_{K,\overline{E}},\overline{\mathbb{F}}_{\ell})$ then this should always admit a filtration in the derived category whose graded pieces are
$R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))[2d_{b}]\otimes^{\mathbb{L}}_{\mathcal{H}(J_{b})}R\Gamma(\mathrm{Ig}^{b},\overline{\mathbb{F}}_{\ell})$
@@
    $\overline{\mathbb{F}}_{\ell}(d_{b})$ is the sheaf on $\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}}$ with trivial Weil group action and $J_{b}(\mathbb{Q}_{p})$ action as defined in [37, Lemma 7.4].
Such a filtration should always exist, but is not currently proven in general. In the case that the Shimura datum $(\mathbf{G},X)$ is PEL of type AC, a modern proof of this result can be found in [37, Theorem 7.1].
@@
We first discuss point (1). One of the key observations underlying Koshikawa’s method was that the cohomology of the space $\mathrm{Sht}(G,b,\mu)_{\infty}$ computes the action of a Hecke operator $T_{\mu}$ corresponding to $\mu$ on $\mathrm{Bun}_{G}$ the moduli stack of $G$-bundles of the Fargues-Fontaine curve. The Hecke operators commute with the action of the excursion algebra on $\mathrm{Bun}_{G}$, and the action of the excursion algebra on a smooth irreducible representation $\rho$, viewed as a sheaf on $\mathrm{Bun}_{G}$, determines the Fargues-Scholze parameter of $\rho$. It follows that $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ as a complex of smooth $J_{b}(\mathbb{Q}_{p})$-modules will have irreducible constituents $\rho$ with Fargues-Scholze parameter $\phi_{\rho}^{\mathrm{FS}}$ equal to $\phi_{\mathfrak{m}}$ as conjugacy classes of parameters. When $\mathbf{G}_{\mathbb{Q}_{p}}=G$ is a product of $\mathrm{GL}_{n}$s, as in Theorem 1.3 (by the assumption that $p$ splits in $F$), it follows from the work of Hansen-Kaletha-Weinstein [31, Theorem 1.0.3] that the Fargues-Scholze correspondence for $J_{b}(\mathbb{Q}_{p})$ with rational coefficients agrees with the semi-simplification of the Harris-Taylor correspondence, where we recall that $J_{b}$ is a product of inner forms of $\mathrm{GL}_{n}$s in this case. In particular, using that $\mathfrak{m}$ is generic, it follows that $\phi_{\rho}^{\mathrm{FS}}=\phi_{\mathfrak{m}}$ must lift to a $\overline{\mathbb{Z}}_{\ell}$ parameter which is also generic in the analogous sense, and the condition of generic implies that the lift cannot come from the semi-simplification of a parameter with non-trivial monodromy. Using this, one can deduce that such a $\rho$ only exists if the group $J_{b}$ is quasi-split. In this particular case ($G$ is a product of $\mathrm{GL}_{n}$s), this can only happen if $b\in B(G,\mu)$ corresponds to the ordinary element (cf. Remark 4.28).
@@
This argument of Koshikawa was formalized and generalized further in work of the first author [24]. In particular, it was noted that, for a general quasi-split $G$ and $\mathfrak{m}$ generic, the cohomology $R\Gamma_{c}(\mathrm{Sht}(G,b,\mu)_{\infty,C}/\underline{K_{p}^{\mathrm{hs}}},\overline{\mathbb{F}}_{\ell}(d_{b}))_{\mathfrak{m}}$ will only be non-trivial if $b\in B(G,\mu)_{\mathrm{un}}:=B(G)_{\mathrm{un}}\cap B(G,\mu)$, where $B(G)_{\mathrm{un}}$ is the set of elements lying in the image of the map $B(T)\rightarrow B(G)$, assuming that the Fargues-Scholze local Langlands correspondence has certain expected properties (Assumption 4.5). These unramified elements will be precisely the elements for which $J_{b}$ is quasi-split. The set $B(G,\mu)_{\mathrm{un}}$ corresponds to Weyl group orbits of weights in the representation $V_{\mu}$ of $\hat{G}$ restricted to $\hat{G}^{\Gamma}$. In particular, if $G$ is split then, since $\mu$ is minuscule, $B(G,\mu)_{\mathrm{un}}$ consists of only one element, corresponding to the unique Weyl group orbit of the highest weight. This is the situation occurring in the previous paragraph. Moreover, the contribution of the cohomology of this shtuka space is easily understood, and the problem completely reduces to controlling the cohomology of $\mathrm{Ig}^{b}$ when $b\in B(G,\mu)_{\mathrm{un}}$ is the $\mu$-ordinary element. However, if $G$ is not split then the restriction of $V_{\mu}$ to $\hat{G}^{\Gamma}$ may have multiple Weyl group orbits of weights. In particular, one needs to control the cohomology groups